# Dual Prototyping with Domain and Class Prototypes for Affective Brain-Computer Interface in Unseen Target Conditions

### v1

@@ III-B Prototype Inference
For domain features, we assume that each domain has a prototype representation, which we refer to as the domain prototype. This prototype represents the key characteristics of that domain, with the distribution of domain features centered around it. For each domain, the domain prototype can be considered the ”centroid” of all its domain features. Similarly, for each category within a domain, we derive class prototypes through prototype inference. The class prototypes capture the essential properties of each class within the domain and serve as the ”centroid” of the class features. Both types of prototypes, domain and class, can be computed as the average value of their respective sample features, denoted as . Specifically, the estimation of domain prototypes for each domain is given by:
@@
where represents the collection of domain features for all samples from the -th subject in the source domain. Here, denotes the number of samples from this subject, is the the domain feature of the -th sample, and is the corresponding domain label for that sample’s domain feature. For data from the same domain, the domain labels are identical. For the class features within a single domain , the class prototype is given as
@@
where represents the set of class features for samples that classified as from the -th subject’s samples. Here, denotes the number of samples classified as in the -th subject’s data. is the class feature of the -th sample, and is the class label for that sample. In summary, for each sample, we will obtain the corresponding domain prototype and class prototypes . Here, represents the number of classes. During training, each subject’s prototypes are calculated based on the features of all their samples and are updated throughout the training process to better capture the feature distribution. In the testing phase, these prototypes are fixed.
@@
After obtaining the domain prototypes and class prototypes, we proceed with prototype inference. For each sample, after feature disentanglement and extraction of the corresponding domain and class features, we first perform domain prototype inference to identify the most suitable domain. This is followed by class prototype inference within the selected domain to determine the class label. Specifically, for the domain feature , we compare its similarity with each class domain prototype using a bilinear transformation as
@@
where is a trainable, randomly initialized bilinear transformation matrix that is not constrained by positive definiteness or symmetry. The model updates the weights of this bilinear matrix through backpropagation, with the purpose of enhancing the feature representation capability for downstream tasks. We compare the similarity between the sample’s domain feature and each domain prototype, as follows:
@@
Here, represents the domain prototype of the -th subject.The most similar domain is determined based on the corresponding to the maximum value in the vector . For each training epoch, once the most similar domain for the sample is identified, the class prototypes for that domain are used to measure the similarity between the sample’s class features and each class prototype. When comparing class features with class prototypes, we use cosine similarity, as follows:
@@
III-C Pairwise Learning
To address this issue and enhance the model’s resistance to label noise, we employ a pairwise learning strategy to replace pointwise learning. Unlike pointwise learning, pairwise learning takes into account the relationships between pairs of samples, capturing their relative associations through pairwise comparisons. The pairwise loss function used is defined as follows.
@@ III-C Pairwise Learning
Here, and are the feature vectors of the class feature of samples and , obtained through prototype inference ( Eq. 9). The symbol represents the dot product operation. The result of falls within the range , representing the similarity between the two feature vectors and . In summary, the objective function for the pairwise learning is defined as follows:
@@
IV Experimental Results
IV-A Dataset and Data Preprocessing
We validate the proposed PL-DCP using the widely recognized public databases SEED [36] and SEED-IV [37]. The SEED dataset includes 15 subjects, each participating in three experimental sessions conducted on different dates, with each session containing 15 trials. During these sessions, video clips were shown to evoke emotional responses (negative, neutral, and positive) while EEG signals were simultaneously recorded. For the SEED-IV dataset, it includes 15 subjects, each participating in three sessions held on different dates, with each session consisting of 24 trials. In this dataset, video clips were used to induce emotions of happiness, sadness, calmness, and fear in the subjects.
The acquired EEG signals undergo preprocessing as follows. First, the EEG signals are downsampled to a 200 Hz sampling rate, and noise is manually removed. The denoised data is then filtered using a bandpass filter with a range of 0.3 Hz to 50 Hz. For each experiment, the signals are segmented using a 1-second window, and differential entropy (DE) features, representing the logarithmic energy spectrum of specific frequency bands, are extracted based on five frequency bands: Delta (1-3 Hz), Theta (4-7 Hz), Alpha (8-12 Hz), Beta (14-30 Hz), and Gamma (31-50 Hz), resulting in 310 features for each EEG segment (5 frequency bands × 62 channels). Finally, a Linear Dynamic System (LDS) is applied to smooth all obtained features, leveraging the temporal dependency of emotional changes to filter out EEG components unrelated to emotions and those contaminated by noise [38]. The EEG preprocessing procedure adheres to the same standards as previous studies to enable fair comparisons with models presented in previous literature.
IV-B Experiment Protocols
@@
IV-C Cross-Subject Single-Session Leave-One-Subject-Out Cross-Validation.
Table II and Table III present the experimental results on the SEED and SEED-IV datasets under cross-subject single-session leave-one-subject-out cross-validation. On the SEED dataset, our proposed PL-DCP model demonstrates a significant performance advantage over both traditional machine learning and non-deep transfer learning methods. Compared to the best-performing non-deep transfer learning method (CORAL), our model shows an improvement of 11.40% (CORAL: 71.48%; PL-DCP: 82.88%). Notably, while other deep transfer learning methods incorporate target domain data during training, our model trains without using any target domain data. Despite this, our method achieves results comparable to, and frequently surpassing, other deep transfer learning approaches that use target domain data, with a 7.46% improvement over DDC and a 5.23% improvement over MS-MDA. On the SEED-IV dataset, PL-DCP also achieves superior results without target domain data in training. Compared to the highest-performing traditional machine learning method, RF (Random Forest: 52.67%), and deep learning methods that utilized target domain data during training (MMD: 59.34%), PL-DCP achieves an accuracy of 65.15% ± 10.34%, outperforming all other methods in both the traditional and deep learning categories. This indicates a substantial improvement in recognition accuracy on the SEED-IV dataset.
@@
IV-D Cross-Subject Cross-Session Leave-One-Subject-Out Cross-Validation.
Compared to cross-subject single-session, cross-subject cross-session not only accounts for variability among subjects but also incorporates differences across sessions. In EEG-based emotion recognition tasks, this evaluation scheme presents the greatest challenge to the model’s effectiveness. As shown in Table IV and Table V, the proposed method achieves a recognition accuracy of 79.34% ± 6.34% on the SEED dataset and 63.16% ± 9.03% on the SEED-IV dataset. Compared to the best results from existing studies (SEED: 78.42% with DANN; SEED-IV: 61.44% with DANN), the proposed method outperforms classic deep transfer learning methods, achieving higher accuracy without utilizing target domain data. These results suggest that the proposed PL-DCP can maintain robust performance independently of target domain data, effectively handling the challenges posed by inter-subject and inter-session variability in EEG-based emotion recognition tasks. This demonstrates the model’s strong validity and generalization capabilities.
@@
| KNN*[39] | 40.06±4.98 | KPCA*[40] | 47.79±7.85 |
| SVM*[41] | 48.36±7.51 | SA*[21] | 40.34±5.85 |
| TCA*[20] | 43.01±7.13 | CORAL*[42] | 50.01±7.93 |
| GFK*[23] | 43.48±6.27 | RF*[43] | 48.16±9.43 |
@@
| Methods | Methods | ||
| DAN* [44] | 60.95±9.34 | DANN*[45] | 61.44±11.66 |
| DCORAL* [46] | 59.96±9.03 | DDC*[24] | 54.76±9.02 |
| PL-DCP | 63.16±09.03 | ||
IV-E Confusion Matrix
To qualitatively assess the performance of the proposed model across different emotion categories, we visualize the confusion matrix and compare the results with classic deep transfer learning methods. As shown in Fig. 3, all models perform best in recognizing positive emotions, while their accuracy declines for negative and neutral emotions. For example, DANN achieves a recognition rate of only 75.96% for neutral emotions. Moreover, the three deep transfer learning methods used for comparison show performance fluctuations across different emotion categories, indicating lower stability. For instance, in DANN, the difference in performance between recognizing positive and neutral emotions is 9.03%. For DAN, the difference between positive and neutral recognition rates is 8.69%. DCORAL exhibits a performance gap of 10.2% between recognizing positive and negative emotions. In contrast, the proposed model maintains consistently high performance across all emotion categories, with recognition rates exceeding 80% for each category. It also demonstrates better stability, even without target domain data during training, with a maximum performance difference of only 4.83% across emotion types. This suggests that the model exhibits enhanced robustness and balanced performance across various emotional states.
V Discussion
V-A Ablation Study
To thoroughly evaluate the model’s performance and assess the impact of each module in the proposed PL-DCP, we conduct an ablation study. The ablation results, based on the SEED dataset under cross-subject single-session leave-one-subject-out cross-validation, are presented in Table VI. (1) Removing the domain discriminator loss. A decrease of 7.53% in model performance is observed. (2) Removing the class discrimination loss. The model’s performance drops from 82.88% to 79.18%, resulting in a decrease of 3.70%. (3) Removing both domain discriminator loss and class discrimination loss. It causes a significant performance decline of 8.39%. This indicates that the combined presence of both domain and class discriminators enhances the extraction of relevant features, substantially improving model recognition performance in the target domain. (4) Removing the bilinear transformation matrix . The bilinear transformation matrix in Eq. 8 contributes to model performance, increasing accuracy by 2.24%. (5) Removing soft regularization . The soft regularizer also improves accuracy, raising it by 1.15%. These results demonstrate the effectiveness of each component within the PL-DCP model and their combined impact on overall performance.
@@ V-A Ablation Study
| w/o domain disc. loss | 75.24/10.52 |
| w/o class disc. loss | 79.18/06.62 |
| w/o domain disc. loss and class disc. loss | 74.49/05.49 |
| w/o the bilinear transformation matrix S | 80.64/04.60 |
| w/o soft regularization R | 81.73/03.31 |
| PL-DCP | 82.88/05.32 |
V-B Visualization of Domain and Class Features
@@
To intuitively understand the extracted domain and class features, we use T-SNE [49] to visualize these features for the respective samples, enabling us a clear observation of how the features and prototypes evolve over training. The visualizations of domain and class features at different training stages are shown in Fig. 4 (a)-(c) and Fig. 4 (d)-(f), respectively. These figures capture the features at the beginning of training (first column), after 50 training epochs (second column), and at the end of training (third column). In the domain feature visualizations (upper row), different colors represent the domain features for each domain, while diamonds in corresponding colors denote the domain prototypes for each domain. The target data are represented as semi-transparent black crosses (×) to avoid excessive overlap with other domain features. Comparing the feature distribution from Fig. 4 (a) to (c), it becomes evident that the domain features for the same subject are closely clustered, forming separate groups with the domain prototypes located at the center of each cluster. This differentiation in domain feature distributions across subjects supports our hypothesis that domain features, derived through feature disentanglement of shallow features, can effectively distinguish between subjects. A similar trend is observed in the class feature visualizations. Here, different colors represent different classes, with the semi-transparent black crosses (×) again indicating the target data. As training progresses, a more defined class boundary among different classes becomes apparent from Fig. 4 (d) to (f), illustrating the model’s ability to learn clear class separability over the course of training. To further illustrate the relationships between domain prototypes and class prototypes (Fig. 5), we analyze both close pairs and distant pairs of domain samples, visualizing their respective representations in the class prototype space. The results reveal that samples closer in the domain prototype space tend to remain closer in the class prototype space, indicating consistency and coherence in the mapping across the two prototype spaces. This observation reinforces the effectiveness of the proposed framework in preserving the intrinsic relationships between samples during dual prototype learning.
V-C Effect of Noisy Labels
We further evaluate the model’s performance under label noise to assess the robustness and noise-resistance capability introduced by pairwise learning. In this process, we randomly replace a proportion (%) of labels in the source domain data with incorrect labels, simulating real-world scenarios where data labels may contain noise. The model is then trained using this noisy source domain data, and its performance is validated on the unseen target domain data. We vary the % values to 5%, 10%, 20%, and 30%, respectively, yielding model accuracies and standard deviations of 81.46% ± 5.54%, 80.32% ± 6.39%, 79.79% ± 6.09%, and 79.01% ± 7.46%. These results indicate that as the label noise rate % gradually increases from 5% to 30%, the model’s performance decreases only slightly, with a steady downward trend resulting in an overall performance drop of just 3.87%. This suggests that the proposed model demonstrates strong robustness against label noise, as its performance does not exhibit significant declines in the presence of noisy labels.
VI Conclusion

### v2

@@ 3.2 Prototype Inference
For domain features, we assume that each domain has a prototype representation, which we refer to as the domain prototype. This prototype represents the key characteristics of that domain, with the distribution of domain features centered around it. As shown in Fig.1(a), For each domain, the domain prototype can be considered the ”centroid” of all its domain features. Similarly, for each class within a domain, we derive class prototypes through prototype inference. As shown in Fig.1(b), the class prototypes capture the essential properties of each class within the domain and serve as the ”centroid” of the class features. Both types of prototypes, domain and class, can be computed as the average value of their respective sample features, denoted as . Specifically, the estimation of domain prototypes for each domain is given by:
@@
where represents the collection of domain features for all samples from the -th subject in the source domain. Here, denotes the number of samples from this subject, is the the domain feature of the -th sample, and is the corresponding domain label for that sample’s domain feature. For data from the same domain, the domain labels are identical. For the class features within a single domain , the class prototype is defined as:
@@
where represents the set of class features for samples that classified as from the -th subject’s samples. Here, denotes the number of samples classified as in the -th subject’s data. is the class feature of the -th sample, and is the class label for that sample. In summary, for each sample, we will obtain the corresponding domain prototype and class prototypes . Here, represents the number of classes. During training, the prototype of each subject is computed based on the features of all its samples. Prototype features are updated iteratively throughout the training process to better capture the feature distribution. Noteworthy, these prototypes are fixed during the testing phase.
@@
After obtaining the domain prototypes and class prototypes, we proceed with prototype inference. For each sample, as shown in Fig.3.(a), we after feature decoupling and extracting the corresponding domain and class features, we first perform domain prototype inference to identify the most suitable domain. Then,as shown in Fig.3.(b), class prototype inference within the selected domain to determine the class label. Specifically, for the domain feature , we compare its similarity with each class domain prototype using a bilinear transformation as:
@@
where is a trainable, randomly initialized bilinear transformation matrix that is not constrained by positive definiteness or symmetry. The model updates the weights of this bilinear matrix through backpropagation, with the purpose of enhancing the feature representation capability for downstream tasks. We compare the similarity between the sample’s domain feature and each domain prototype, defined as:
@@
here, represents the domain prototype of the -th subject.The most similar domain is determined based on the corresponding to the maximum value in the vector . For each training epoch, once the most similar domain for the sample is identified, the class prototypes for that domain are used to measure the similarity between the sample’s class features and each class prototype. When comparing class features with class prototypes, we use cosine similarity, defined as:
@@
3.3 Pairwise Learning
To solve the label noise problem and enhance the model’s resistance to label noise, we employ a pairwise learning strategy to replace pointwise learning. Unlike pointwise learning, pairwise learning takes into account the relationships between pairs of samples, capturing their relative associations through pairwise comparisons. The pairwise loss function used is defined as follows.
@@ 3.3 Pairwise Learning
Here, and are the feature vectors of the class feature of samples and , obtained through prototype inference (Eq.10). The symbol () represents the dot product operation. The result of falls within the range [0,1], representing the similarity between the two feature vectors and . In summary, the objective function for the pairwise learning is defined as follows:
@@
4 Experimental Results
4.1 Dataset and Data Preprocessing
We validate the proposed PL-DCP using the widely recognized public databases SEED [46], SEED-IV [47] and SEED-V[48]. The SEED dataset includes 15 subjects, each participating in three experimental sessions conducted on different dates, with each session containing 15 trials. During these sessions, video clips were shown to evoke emotional responses (negative, neutral, and positive) while EEG signals were simultaneously recorded. For the SEED-IV dataset, it includes 15 subjects, each participating in three sessions held on different dates, with each session consisting of 24 trials. In this dataset, video clips were used to induce emotions of happiness, sadness, calmness, and fear in the subjects. In the SEED-V database, a total of 16 subjects were participated, with each completing three sessions on different dates. Each session comprised 15 trials and include five emotions: happiness, neutral, sadness, disgust and fear.
The acquired EEG signals undergo preprocessing as follows. First, the EEG signals are downsampled to a 200 Hz sampling rate, and noise is manually removed, such as electromyography (EMG) and electrooculography (EOG). The denoised data is then filtered using a bandpass filter with a range of 0.3 Hz to 50 Hz. For each experiment, the signals are segmented using a 1-second window, and differential entropy (DE) features, representing the logarithmic energy spectrum of specific frequency bands, are extracted based on five frequency bands: Delta (1-3 Hz), Theta (4-7 Hz), Alpha (8-12Hz), Beta (14-30Hz), and Gamma (31-50Hz), resulting in 310 features for each EEG segment (5 frequency bands × 62 channels). Finally, a Linear Dynamic System (LDS) is applied to smooth all obtained features, leveraging the temporal dependency of emotional changes to filter out EEG components unrelated to emotions and those contaminated by noise.[49] The EEG preprocessing procedure adheres to the same standards as previous studies to enable fair comparisons with models presented in previous literature. All models are executed under the following configuration: NVIDIA GeForce RTX 3090, CUDA=11.6, PyTorch =1.12.1.
4.2 Experiment Protocols
@@
| Methods | Methods | ||
@@
| Methods | Methods | ||
@@
| KNN*[50] | 35.73±07.98 | KPCA*[51] | 35.47±09.39 |
| SVM*[52] | 53.14±10.10 | SA*[29] | 36.06±11.55 |
| TCA*[28] | 37.57±13.47 | CORAL*[53] | 55.18±07.42 |
| GFK*[31] | 38.32±10.11 | RF*[54] | 42.29±16.02 |
| Deep learning methods | |||
| DAN*[55] | 59.36±16.83 | DANN*[56] | 56.28±16.25 |
| DCORAL*[57] | 56.26±14.56 | DDC*[58] | 56.54±18.35 |
| PL-DCP | 61.29±09.61 | ||
@@
| KNN*[50] | 40.06±04.98 | KPCA*[51] | 47.79±07.85 |
| SVM*[52] | 48.36±07.51 | SA*[29] | 40.34±05.85 |
| TCA*[28] | 43.01±07.13 | CORAL*[53] | 50.01±07.93 |
| GFK*[31] | 43.48±06.27 | RF*[54] | 48.16±09.43 |
| Deep learning methods | |||
| DAN*[55] | 60.95±09.34 | DANN*[56] | 61.44±11.66 |
| DCORAL*[57] | 59.96±09.03 | DDC*[58] | 54.76±09.02 |
| PL-DCP | 63.16±09.03 | ||
@@
| Traditional machine learning methods | |||
| KNN*[50] | 35.28±07.57 | KPCA*[51] | 39.68±11.28 |
| SVM*[52] | 41.20±10.76 | SA*[29] | 31.87±09.87 |
| TCA*[28] | 37.68±08.40 | CORAL*[53] | 54.68±07.44 |
| GFK*[31] | 37.89±09.84 | RF*[54] | 43.63±11.38 |
@@
| DAN*[55] | 54.27±10.42 | DANN*[56] | 52.83±13.90 |
| DCORAL*[57] | 52.23±12.76 | DDC*[58] | 43.89±11.84 |
| PL-DCP | 57.53±09.72 | ||
4.3 Cross-Subject Single-Session Leave-One-Subject-Out Cross-Validation.
The experimental results of PL-DCP on the SEED database are shown in Tab.2, our proposed PL-DCP model demonstrates a significant performance advantage over both traditional machine learning and non-deep transfer learning methods. Compared to the best-performing non-deep transfer learning method (CORAL), our model shows an improvement of 11.40% (CORAL: 71.48%; PL-DCP: 82.88%). Notably, while other deep transfer learning methods incorporate target domain data during training, our model trains without using any target domain data. Despite this, our method achieves results comparable to, and frequently surpassing, other deep transfer learning approaches that use target domain data, with a 7.46% improvement over DDC and a 5.23% improvement over MS-MDA. The results on the SEED-IV dataset are shown in Tab.3, PL-DCP also achieves superior results without target domain data in training. Compared to the highest-performing traditional machine learning method (RF: 52.67%), and deep learning methods that utilized target domain data during training (MMD: 59.34%), PL-DCP achieves an accuracy of 65.15% ± 10.34%, outperforming all other methods in both the traditional and deep learning methods. The results on the SEED-V database are shown in Tab.4. The results of PL-DCP are better than those of traditional machine learning methods. The PL-DCP model achieves an accuracy of 61.29%, which is 1.93% higher than the suboptimal method (DAN: 59.36%).
All these results prove that PL-DCP achieves superior performance even without relying on target domain data, and has strong robustness and generalization ability.
4.4 Cross-Subject Cross-Session Leave-One-Subject-Out Cross-Validation.
Compared to cross-subject single-session, cross-subject cross-session not only accounts for variability among subjects but also incorporates differences across sessions. In EEG-based emotion recognition tasks, this evaluation scheme presents the greatest challenge to the model’s effectiveness. The test results of the PL-DCP model in the SEED database are shown in Tab.5. The PL-DCP model achieves an accuracy of 79.34%, which is significantly higher than that of traditional machine learning methods and deep transfer learning methods. Compared with the suboptimal model (DANN: 78.24%), the performance of the PL-DCP model is improved by 0.92%. As shown in Tab.6, the PL-DCP model achieves the best emotion recognition performance in the SEED-IV database, with an accuracy of 63.16%, which is significantly improved compared with traditional machine learning methods and deep learning methods. Compared with the suboptimal model (DANN: 61.44%), the performance of PL-DCP is improved by 1.72%. The test results of the PL-DCP model on the SEED-V database are shown in Tab.7, the classification performance of 57.53% ± 09.72% was achieved in the five EEG emotion categories of the SEED-V database, compared with the suboptimal performance (CORAL: 54.68%), the accuracy is improved by 2.85%. It is worth noting that although the target domain data is completely unknown during training, the PL-DCP model still achieves slightly better performance than the deep transfer learning method that requires both source and target data.
All these results suggest that the proposed PL-DCP can maintain robust performance independently of target domain data, effectively handling the challenges posed by inter-subject and inter-session variability in EEG-based emotion recognition tasks, which demonstrates the model’s strong validity and generalization capabilities.
4.5 Confusion Matrix
The overall accuracy reflects the global performance of the model, but in the practical application of emotional brain-computer interface, the misjudgment of a specific emotion class may cause completely different interaction consequences. In order to qualitatively evaluate the performance of the proposed model on different sentiment categories, we visualize the results of leave-one-out cross-validation on the SEED, SEED-IV and SEED-V datasets using confusion matrices, and compare the results with deep transfer learning methods. To measure the recognition performance of the model for different emotions.
On the SEED dataset, as shown in Fig.4.(a)(d), all models have the best performance in recognizing positive emotion, which are 81.86%, 80.16%, 81.50%, 78.85%, respectively. In contrast, the recognition accuracy of negative and neutral emotions is slightly lower. In addition, the performance difference between positive and neutral emotion of DANN model is 8.69%, the performance difference between positive and neutral emotion of DAN model is 8.77%, and the performance difference between positive and negative emotion of DCORAL model is 10.4%. However, the maximum performance difference of PL-DCP between emotion categories is only 4.09%. Obviously, PL-DCP showed more robust and balanced performance, while other deep transfer learning methods showed lower stability with fluctuating performance on different emotion categories.
On the SEED-IV dataset, as shown in Fig.4.(e)(h), On the SEED-IV dataset, the recognition accuracy of all models for neutral emotion is slightly higher, reaching 76.53%, 65.96%, 64.12% and 61.35%, respectively, while the recognition accuracy for happy emotion is slightly lower. In addition, on this dataset, each model is easy to confuse sadness and fear.
On the SEED-V dataset, as shown in Fig.5.(a)(d), all the models have slightly lower recognition accuracy performance for happy emotions, while they have higher recognition accuracy for neutral, sad, and fear emotions. All models easily confuse happy emotions with fear and neutral emotions. Compared with other models, the PL-DCP model has the best performance in recognizing neutral emotions. Overall, our proposed PL-DCP model has superior emotion recognition performance and stability.
@@ 4.5 Confusion Matrix
| w/o domain prototype | 74.67±07.62 |
| w/o domain disc. loss | 75.24±10.52 |
| w/o class disc. loss | 79.18±06.62 |
| w/o domain disc. loss and class disc. loss | 74.49±05.49 |
| w/o pairwise learning | 77.61±09.20 |
| w/o the bilinear transformation matrix S | 80.64±04.60 |
| w/o soft regularization R | 81.73±03.31 |
| PL-DCP | 82.88±05.32 |
5 Discussion and Conclusion
5.1 Ablation Study
To comprehensively evaluate the performance of the model and evaluate the impact of each module in the proposed PL-DCP model, we performed ablation experiments. Ablation results under cross-subject single-session leave-one-subject-out cross-validation based on the SEED dataset are shown in Tab.8. After removing domain prototype, it is observed that the performance of the model decreases by 8.21% when only the class prototype is used, indicating that the domain and class dual prototype representation method has performance advantages compared with the single class prototype representation method. After removing the domain discriminator loss, we observe a 7.53% decrease in model performance. After removing the class discriminator loss, the model performance decreases from 82.88% to 79.18%, with a decrease of 3.70%. After removing domain discriminator loss and class discriminator loss, the model performance decreases significantly by 8.39%, which indicates that the interaction of domain discriminator and class discriminator helps to extract relevant features and significantly improves the recognition performance for the target domain. After removing the pairwise learning strategy and replacing it with the pointwise learning strategy, the model performance decreases by 5.27%, indicating that the paired learning strategy enhances the recognition performance of the model by capturing the relationship between sample pairs. Removing bilinear transformation matrix , the model performance decreases by 2.24%, indicating that the bilinear transformation matrix contributes to the model performance in the formula Eq. 8. After removing the soft regularization term , the model performance decreases by 1.15%.
Overall, All these results demonstrate the effectiveness of the individual components in the PL-DCP model and their combined impact on the overall performance.
5.2 Visualization of Domain and Class Features
To intuitively understand the extracted domain and class features, we use T-SNE [62] to visualize these features for the respective samples, enabling us a clear observation of how the features and prototypes evolve over training. The visualizations of domain and class features at different training stages are shown in Fig.6.(a)(c) and (d)(f), respectively. These figures capture the features at the beginning of training (first column), after 50 training epochs (second column), and at the end of training (third column).
@@ 5.2 Visualization of Domain and Class Features
In the visualization of domain features from (a)(c), different colors represent the domain features for each domain, while diamonds () in corresponding colors denote the domain prototypes for each domain. The target data are represented as semi-transparent black crosses (×) to avoid excessive overlap with other domain features. Comparing the feature distribution from Fig.6.(a)(c), it is evident that the domain features of the same subject are clustered more and more closely, forming separate groups with the domain prototypes () located at the center of each cluster. This differentiation in domain feature distributions across subjects supports our hypothesis that domain features, derived through feature disentanglement of shallow features, can effectively distinguish between subjects. A similar trend is observed in the class feature visualization as shown in Fig.6.(d)(f). Here, different colors represent different classes, with the semi-transparent black crosses (×) again indicating the target data. As the training progresses, clearer class boundaries between different categories become more and more obvious, indicating that the model learns a strong feature alignment ability based on class prototypes during training.
To further illustrate the relationships between domain prototypes and class prototypes, we analyze both close pairs and distant pairs of domain samples. As shown in Fig.7.(a)(b), we visualizing their respective representations in the class prototype space.(a) shows that the domain features of the subject pairs are relatively close, with a high degree of overlap, and the class feature distributions are therefore relatively close. (b) shows that the domain features of the subject pairs are relatively far away, with a low degree of overlap, and the class feature distributions are therefore relatively far away. The results reveal that Samples that are close to each other in the domain prototype space also tend to remain close in the class prototype space, indicating consistency and coherence in the mapping across the two prototype spaces. This observation reinforces the effectiveness of the proposed framework in preserving the intrinsic relationships between samples during dual prototype learning.
5.3 Effect of Noisy Labels
We further evaluate the performance of the model under label noise to evaluate the robustness and noise immunity introduced by pairwise learning. During the experiment, the labels of % (5%, 10%, 20% and 30%) of the source domain data are randomly replaced with random error labels, which simulates the situation that the data labels contain noise in the real scene. We tested the PL-DCP model using cross-subject single-session leave-one-out cross validation on the SEED database, and compared it on pointwise learning and pairwise learning strategies.
The results are shown in Tab.9. When using the pointwise learning strategy, The accuracy of the model in different proportions of label noise was 77.61%, 76.76%, 73.49%, 69.17%, 66.53%. The performance of the model decreases rapidly with the increase of , and the performance of the model decreases by 11.08% when the label noise rate reaches 30%. When using pairwise learning strategy, the accuracy rates are 81.46%, 80.32%, 79.79% and 79.01%, respectively. As the label noise rate gradually increases from 5% to 30%, the overall performance of the model only decreases by 3.87%. The results show that the performance of the PL-DCP model only slightly decreases in the label noise environment, showing superior performance, and the introduction of pairwise learning strategy shows strong robustness under label noise.
| Noisy Ratio | Pointwise Learning | Pairwise Learning |
| 0% | 77.6109.20 | 82.8805.32 |
| 5% | 76.7608.64 | 81.4605.54 |
| 10% | 73.4907.96 | 80.3206.39 |
| 20% | 69.1708.31 | 79.7906.09 |
| 30% | 66.5306.63 | 79.0107.46 |
5.4 Conclusion